Kenisha sells athletic shoes part time at a department store. She can earn either $500 per month plus a 4% commission on her total sales, or $400 per month plus a 5% commission on total sales. What is the total price of the athletic shoes Kenisha needs to sell to earn the same income from each pay scale.
I only have a clear idea on the system of equations.
500+0.04c
400+0.05c

Answer by rfer(16322)

500+0.04c=400+0.05c
500-400=0.05c-0.04c
100=0.01c
100/0.01=c
c=$10,000.00